OpenStudy (anonymous):

If x>0 and y>0 and x^9 = 4 and x^7 = 9/y^2, which of the following is an expression for the value of x in terms of y? A) 4y/9 B) 2y/3 C) 3y^2/2 D) 6y or E) 36y^2 ?

OpenStudy (anonymous):

x^9 / x^7 = x^2 = 4 / 9/y^2 x^2 = 4y^2 / 9 x = 2y / 3 answer is B

OpenStudy (anonymous):

thanx